find the solution of the following D,E                                         dy/dx=sin(x-y)               ,,ineed your help guys

 

Grade:12th pass

1 Answers

Arun
5 years ago
Put x-y = u
Then
1 - dy/dx = du/dx
dy/dx = 1 - du/dx
1 - du/dx = sinu
du/dx = 1 - sin u = (sinu/2 + cos u/2)²
Now you can solve